OpenStudy (calculusxy):

A sunflower is 3 inches tall at week 0 and grows 2 inches each week. Does the function `f(n) = f(n - 1) + 2 where f(0) = 3` represent the height, f(n), in "n" weeks?

OpenStudy (calculusxy):

@mathstudent55 @mathmale @agent0smith

OpenStudy (calculusxy):

The answer key actually says that it does. But I don't know why.

OpenStudy (faiqraees):

It should be this General formula for arithmetric progression = a +(n-1)d \[\large\rm F_n = 3+ 2n\]

OpenStudy (agent0smith):

The given formula should work. f(n-1) represents the height last week, and adding 2 matches the description.

OpenStudy (faiqraees):

@agent0smith What's the value of f(n-1) when n = 0 and how's f(0)=3

OpenStudy (agent0smith):

f(0) = 3 is a given. You don't use the function to find it. The next week would be n=1. The formula is correct.

OpenStudy (faiqraees):

How's the formula satisfying f(0) = 3

OpenStudy (agent0smith):

Did you not read what I just wrote? f(0) = 3 is a given initial value.

OpenStudy (faiqraees):

like mine satisfies f(0) = 3 +2n f(0) = 3+2(0) f(0) = 3

OpenStudy (agent0smith):

f(1) = f(0) + 2 f(2) = f(1) + 2 and so on. It's a recursive formula. It is correct. It's equivalent to yours.